Solve the linear equation given the following:
Parallel to 5x-y=3 and passing through the point (3,-1)

Answer by Cromlix(4381)

Rearrange 5x - y = 3
into y = mx + c form.
y = 5x - 3
Lines that are parallel to one
another have the same slope (m)
y = 5x - 3
has m = 5
Using the equation:
y - b = m(x - a)with (3,-1)
With a = 3, b = -1 and m = 5
y - (-1) = 5(x - 3)
y + 1 = 5x - 15
y = 5x -15 -1
y = 5x - 16.
Hope this helps.
:-)